Herschel Status Report - October 2010
The second Direct Liquid Contents Measurement is planned to be performed on 2 November 2010, approximately one year after the first such measurement on 25 November 2009 (see status report from December 2009). This measurement assesses the remaining amount of liquid helium (LHe) for the spacecraft's active cooling system, which directly determines Herschel's lifetime for scientific observations. The measurement should allow confirming whether or not the average LHe flow rate accurately reflects the predictions of the thermal model that underlies the life time calculations.
Tests of a new pointing mode ("on-board gyro propagation") are in progress (11 and 21 October 2010) to investigate whether the lower limit of the PACS spectroscopic sensitivity can be improved. The main challenge is one of reproducing, on a timescale of 15 minutes and with sub-arcsec accuracy, a repeated pattern of small pointing offsets. This reproducibility would allow co-adding (stacking) of the detector frames taken on-source and off-source to be done blindly; meaning without having to apply small corrections for positional inaccuracy (which translate into small shifts in line frequency in the slit spectrum).
Apart from a single event upset with SPIRE, all three instruments continue to operate nominally.
A PACS on-board software modification has been uploaded successfully to mitigate the effects of an anomaly affecting the spectroscopy mode (spontaneous changes in some voltage reference settings - see previous reports from July/August and September). PACS photometry mode is not affected and continues nominally. Test observations in the spectroscopy mode have resumed and since the introduction of the update the anomaly has not recurred.
Practically all of the observing time coming free by temporarily suspending observations in PACS spectroscopy mode was used for additional HIFI observations. Somewhat fortunately, this was in a period of time during which the Galactic Centre and Orion regions were visible which, due to their high IR luminosities, are favourite locations for observations at the very high spectral resolution provided by HIFI. The average fraction to which HIFI observations from Key Programmes have been completed has now fully caught up with the other two instruments, making up for the delay that resulted from the late start of HIFI routine science operations.
In response to the first Open Time (OT1) Call for proposals, 576 proposals had been received. The Herschel Observing Time Allocation Committee (HOTAC) has prepared its recommendations at a meeting that was held from 11 to 14 October 2010, at ESAC. The approved OT1 proposals are planned to be announced on 1 November, following endorsement of the committee's recommendations by the ESA Director of Science and Robotic Exploration. The successful Open Time proposers can subsequently finalise their deliveries of schedulable observations from the beginning of November onward, taking comments from the HOTAC into account as appropriate. The observations will then be available for scheduling, which could lead to the first non Key Programme OT observations with Herschel being included in the mission timeline, for scheduled execution in early 2011.
During the reporting period, mission operations have been conducted with the support of the ESA New Norcia ground station. Observational data stored on-board Herschel was received on ground during communication passes lasting approximately three hours.
The ground segment is operating nominally. Data products are generated routinely and ingested into the Herschel Science Archive (HSA).
- 1 November 2010: Announcement of approved OT1 proposals
- 2 November 2010: second Direct Liquid Contents Measurement